A 6.0 ft person weighs 161 lb. Express this person's height in meters and weight in kilograms. (1 lb = 453.6 g; 1 m = 3.28 ft.)
Answer #1

Answer:

Given height of a person=6 ft

We know that 1m=3.28 ft, that means 1 ft=(1/3.28)m=0.305 m.

Therefore 6ft=? m

Height of person= (0.305 m x 6 ft)/(1 ft)=1.8288 m ~ 1.83 m.

Weight of a person=161 lb.

Since 1 lb=453.6 g=0.4536 kg (1 kg=1000 g)

Therefore 161 lb=? Kg

Weight of a person=(161 lb x 0.4536 Kg)/(1 lb)=73.0296 Kg ~ 73.03 Kg.

Therefore the person height ~ 1.83 m and the person weight ~ 73.03 Kg.

Please let me know if you have any doubt. Thanks.